Start learning →

Indian Navy Pilot At Least One — practice questions

Explore 1 free At Least One questions in Indian Navy Pilot Mathematics. Open each question or exercise to read its content, reference answers and available explanation.

Practise At Least One with Quill →

At Least One questions and exercises with answers

Showing questions 1–1 of 1.

Worked practice questions

Ten eggs drawn with replacement from a lot of 10% defective. Probability that **at least one** is defective:

  1. $1 - (9/10)^{10}$
  2. $(1/10)^{10}$
  3. $10 \cdot (1/10)$
  4. $10!$
Show answer and solution

Correct answer: A. $1 - (9/10)^{10}$

$P(X \geq 1) = 1 - P(X = 0) = 1 - {}^{10}C_0 (1/10)^0 (9/10)^{10} = 1 - (9/10)^{10} \approx 1 - 0.349 = 0.651$.

Open this question →

How to practise

  1. Choose a At Least One question and try to work out your answer.
  2. Open the question to check the correct answer and read its worked solution.
  3. Continue in the app for bite-sized lessons, explanations and review with Quill. You can start without creating an account.

More Mathematics topics